32=y-8- 13y
what does y equal?

Answer by checkley77(12844)

32=y-8- 13y
-Y+13Y=-8-32
12Y=-40
Y=-40/12
Y=-10/3 ANS.
PROOF:
32=-10/3-8-13*-10/3
32=-10/3-8+130/3
32=-10/3-24/3+130/3
32=-34/3+130/3
32=96/3
32=32